• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 00:01:01 on October10,2000

    Entry number 49964

    keyword=Swing Shift Summary

    16:00---Start swing shift.
    Continue taking production data at kin PY-2C. Need to take one
    more hour's data for this kin.
    16:09---STop run 1660 for MCC to tune for higher beam.
    16:31---Beam back to HallA. 89.5uA.
    17:06---Jeff adjusted the trigger to setting 10. Take one more hour's data
    at this new trigger.
    18:06---End run 1663. Change to carbon target for pointing
    18:23---Prestart of run 1664 failed. EB1 problem. Kcoda, restart runcontrol.
    Roc14 problem. Kcoda, reset Roc 14.
    19:11---Change Right arm's angle to 45.9 and momentum to 1249.3
    Change to 4He target.
    19:47---Kinamatics change finished.
    20:05---Finish link run data taking, Change Left Arm's momentum to
    Ask MCC to increase current to 90uA.
    20:27---Start production dayta taking at kin PY-2B*.
    21:41---Change to carbon target.
    21:55---Beam position shifted from (0,0). FB off. Ask MCC to turn it on
    and center beam to (0,0). MCC is working on the fast feed back problem.
    22:25---Restart pointing data taking
    23:05---Change Right Arm momentun to 1.10460Gev, angle to 51.50
    23:28---Finish kinematics change
    23:35---Start run 1675, kin=Link B/A.
    23:56---Change Left Arm's momentum to 2.633Gev.

    Run# Time Nevts Comment
    1659 1.2M Continue from last shift, kin=PY-2C
    1660 16:04 0.14M
    1661 16:31 1.2M I=90uA
    1662 17:03 1.2M Change trigger to setting 10
    1663 17:35 1.2M
    1664 18:23 Carbon ponting, raster off
    1665 18:32 1.2M Carbon, raster off
    1666 18:52 0.343M 10 cm dummy, raster on
    1667 19:51 1.2M He4 target, Link run, new right arm setting
    1668 20:27 1.2M Production run kin=PY-2B*
    1669 20:49 1.2M
    1670 21:12 1.2M.
    1671 21:47 0.26M Carbon pointing, raster off, fast feedback off
    1672 22:14 0.6M Carbon pointing, raster off
    1673 22:25 Carbon pointing, junk run
    1674 10cm dummy, raster on
    ------------------------------------------------- 1675 23:34 0.95M He4 target, Link B/A, ROCD 14 problem forced end run
    1676 23:47 0.55M Link B/A